OpenStudy (anonymous):

Vx is the velocity of a particle moving along the x-axis as shown. If x = 2.0 m at t = 1.0 s, what is the position of the particle at t = 6.0 s?

OpenStudy (anonymous):

here is the graph

OpenStudy (anonymous):

These are the answers they give me +2.0 m 6.0 m −1.0 m −2.0 m +1.0 m

OpenStudy (ash2326):

find the area under the curve it'll give the total displacement till t= 1 sec , the area under the curve is =(2-0)*1+1/2 (4-2)*(2-1)=3 units it's given the particle was on x=2 units at t=1 so so it was on x=-1 at the starting Let's find first the area of graph above x axis= 1/2 *(2-0)*(4-0)=4 units in +ve x -axis direction area below the x -axis = 1/2 (6-2)(-2-0)=-4 units so the total displacement will be -4+4=0 so it'll be at it's original position x=-1 option 3

OpenStudy (anonymous):

thanks a bunch it worked

OpenStudy (anonymous):

2+(5*vx)

OpenStudy (anonymous):

please could any one tell me why x=-1